in a street of 40 house 30 receive local newspaper what fraction of the houses get the paper
1
Expert's answer
2020-05-01T17:50:17-0400

fraction of the houses get the paper = number of the houses get the paper / overall number of houses = 30/40 = 3/4 = 0.75

Answer: 3/4 or 0.75